OpenStudy (anonymous):

Let f(x) = twenty-four over the quantity of 3 x minus two. Find f(-2). -6 -3 3 6 can someone explain how to do this?

OpenStudy (kkutie7):

f(x)=\[\frac{ 24 }{ 3x-2}\]

OpenStudy (kkutie7):

plug in -2 into x.

OpenStudy (kkutie7):

\[\frac{ 24 }{ 3(-2)-2 }\] \[\frac{ 24 }{ -6-2 }\] \[\frac{ 24 }{ -8 }\] -3

OpenStudy (anonymous):

thanks, my lesson for virtual school has way more complicated explanations on how to do this, and its been a good 5-6 months since i've done any geometry or algebra :D appreciate the help

OpenStudy (kkutie7):

Yeah any time =)